Quadratic Equation
Z^2-4z+3=0
I worked it this way but can get the right answer can you tell me what I am doing wrong?
Z^2 - 4z + 3 = 0
a=Z
b=-4
c=3
z = 4+- Ѵ4^2-4(z) (3) over 2(z)
z=4+4+- Ѵ4 over 2
answer in back of book 1, 3

Answer by scott8148(6628)

there may be a typo involved; the z's should probably be the same (the variable)

z^2 - 4z + 3 = 0

factoring ___ (z - 1)(z - 3) = 0

z - 1 = 0 ___ z = 1

z - 3 = 0 ___ z = 3